Find x if e^1.7x = 3, and you are given ln3 = 1.0986.
Answer by eperette(173)

Find x if e^1.7x = 3, and you are given ln3 = 1.0986.
Note: e and ln are inverses of each other:
ln (e^1.7x) = ln 3
1.7x = ln3
1.7x/1.7 = ln3/1.7
x = 1.0986/1.7
x = 0.6462